OpenStudy (brucebaner):

A yearbook printer charges based on the number of pages printed. Here is a table that shows the cost of some recent yearbooks: 07_08_LQ_4.gif Which equation gives the cost, y, in terms of the number of pages, x? A. y = 0.012x + 49.985 B. y = 83.3x + 0.65 C. y = 0.012x + 0.65 D. y = 0.012x + 1.25

OpenStudy (brucebaner):

@AriPotta

OpenStudy (brucebaner):

can you help me @AriPotta

OpenStudy (brucebaner):

help me @iambatman

OpenStudy (brucebaner):

i think its D

OpenStudy (aripotta):

take two points, find the slope. then take one point and the slope, and put them into point-slope form: y - y1 = m(x - x1). then convert that into slope-intercept form

OpenStudy (brucebaner):

so em i Right or not is it D

OpenStudy (brucebaner):

@iambatman

OpenStudy (aripotta):

it's not d

OpenStudy (aripotta):

can you show your work?

OpenStudy (brucebaner):

first find found the slope 1.25−1.8550−100=0.012

OpenStudy (brucebaner):

and when i use any point to find the y intercept use point (50,1.25) and we get y = 0.012x+b becomes 1.25=0.012(50)+b so b = .65 and you are CORRECT✓

OpenStudy (aripotta):

so the answer is c, not d :)

OpenStudy (brucebaner):

sorry about what